Can u please integrate :- 3 cos x +4/ sin^2 x....I am unable to solve this

Grade:11

1 Answers

Clement Yung
6 years ago
\int 3\cos (x) + \frac{4}{\sin ^2(x)} dx
= \int 3\cos (x) + 4\csc ^2(x) dx
= 3\sin (x) - 4\cot (x) +C